OpenStudy (anonymous):

How would I solve this equation: e^3x=10

OpenStudy (anonymous):

\[e^{3x}=10\\ 3x=\ln(10)\] is the first step

OpenStudy (michele_laino):

another way to explain, may be this, by definition of logarithm we can write: \[10=e ^{\ln 10}\] so your equation can be rewritten as below: \[e ^{3x}=e ^{\ln(10)}\] now equating the exponents each other, we can write: \[3x=\ln(10)\]

OpenStudy (michele_laino):

subsequently, please divide both sides of that equation, by three, and you will get: \[x=\frac{ \ln 10 }{ 3 }\]

OpenStudy (michele_laino):

please note that: \[\ln 10=\log _{e}10\]